# The Nambu-Goto string as higher-derivative Liouville theory

### v1

@@ V Relation to minimal models
and $\gamma_{0}=1-b^{-2}$ is as in Eq. (1) with the KPZ barriers shifted to $d_{\pm}$ given by (24b).
The values of $d_{\pm}$ depend on $G$ which has to lie in the interval $[-1,0]$ for the the action (10) to be stable. Then $b^{-2}$ is real for $d<d_{-}$ which increases from 1 at $G=0$ to 19 at $G=-1$. For $G=-1/3$ we have $d_{-}=15-4\sqrt{6}\approx 5.2>4$ as is already annonced in Introduction, so $\gamma_{0}$ is real in $d=4$. Remarkably, the value $G=-1/3$ is associated in $d=4$ with the $p=3$, $q=p+1=4$ unitary minimal model as it will be momentarily discussed.
@@
where $b$ is given by Eq. (24a). This $c$ is the central charge of the Virasoro algebra and not to be confused with the central charge $c^{(\varphi)}=26-d$ of $\varphi$ (or the sum of those of $\varphi$ and $\lambda$’s for the Nambu-Goto string).
The relation with the minimal models is established by choosing
$c=25+6\frac{(p-q)^{2}}{pq}.$ (28)
Then Eqs. (24a), (27) fix
$G=\frac{(1-d-6\frac{(p-q)^{2}}{pq})q}{6(q+p)},~{}~{}1-6\frac{(p-q)^{2}}{pq}\leq d\leq 19-6\frac{p}{q}$ (29)
with coprime $q>p$. In the usual Liouville theory where $G=0$ Eq. (29) would imply $d=26-c=1-6\frac{(p-q)^{2}}{pq}$ for the central charge of matter, but $d$ is a free parameter for $G\neq 0$. The inequalities in (29) guarantee $10\geq G\geq-$ 1 as is necessary for stability. Contrary to the Liouville theory now Kac’s $c\neq 26-d$.
Given (29) we finally find from Eq. (24a)
@@
that applies for $d>25-6{\frac{(p+q)^{2}}{p^{2}}}$. The perturbative branch is as in the usual Liouville theory, but the second branch is no longer $p\leftrightarrow q$ interchangeable with it.
@@
VI Final remarks
While our critical indices are as for the perturbative branch in the usual Liouville theory, the domain of applicability is now broader complimenting applications of the Liouville action in Condensed Matter Theory [24, 25].
The indication that conformal symmetry of the Nambu-Goto string and the Ising model on a dynamical lattice are both described by the (4,3) minimal model sounds fascinating and deserving future studies. In particular, direct QFT calculations repeating Salieri’s check in [18] are expected to be done.
Equation (4) represents a “massive” generalization of the Liouville action associated with the conformal anomaly. I expect it may have wide applications and help to make an insight in quantum gravity as the usual Liouville action did [26, 27].
Acknowledgement:
During the work on this project I benefited useful comments from my colleagues. I would like to thank Jan Ambjørn, Andrei Barvinsky, Keshav Dasgupta, Vladimir Kazakov, Charlotte Kristjansen, Alexei Morozov, Gordon Semenoff, Arkady Tseytlin, Dmitry Vasiliev, Mikhail Vasiliev, Alexander Zamolodchikov, Konstantin Zarembo for these. This work was carried out within the research program of NRC “Kurchatov Institute”.

### v3

@@ Introduction
and $\gamma_{0}=1-b^{-2}$ ($\gamma_{h}=1+(h-1)b^{2}$ for surfaces of genus $h$) is as in Eq. (1) with the KPZ barriers shifted to $d_{\pm}$ given by (31b).
The values of $d_{\pm}$ depend on $G$ which has to lie in the interval $[-1,0]$ for the the action (10) to be stable. Then $b^{-2}$ is real for $d<d_{-}$ which increases from 1 at $G=0$ to 19 at $G=-1$. For $G=-1/3$ we have $d_{-}=15-4\sqrt{6}\approx 5.2>4$ as is already announced in the Introduction, so $\gamma_{0}$ is real in $d=4$. Remarkably, the value $G=-1/3$ is associated in $d=4$ with the $p=3$, $q=p+1=4$ unitary minimal model as it will be momentarily discussed.
@@
where $b$ is given by Eq. (31). This $c$ is the central charge of the Virasoro algebra and not to be confused with the central charge $c^{(\varphi)}=26-d$ of $\varphi$ (or $\varphi$ plus $\lambda$’s).
The minimal models are obtained by choosing
$c=25+6\frac{(p-q)^{2}}{pq}$ (35)
with coprime $q>p\geq 2$. Then Eqs. (31), (34) fix
$G=\frac{(1-d-6\frac{(p-q)^{2}}{pq})q}{6(q+p)},~{}~{}1-6\frac{(p-q)^{2}}{pq}\leq d\leq 19-6\frac{p}{q}.$ (36)
In the usual Liouville theory where $G=0$ Eq. (36) implies $d=26-c=1-6\frac{(p-q)^{2}}{pq}$ for the central charge of matter, but $d$ is a free parameter for $G\neq 0$. Contrary to the Liouville theory now $c\neq 26-d$. The inequalities in (36) guarantee $10\geq G\geq-$ 1 as is necessary for stability.
Given (36) we finally find from Eq. (31)
@@
that applies for $d>25-6{\frac{(p+q)^{2}}{p^{2}}}\geq 2$. The perturbative branch is as in the usual Liouville theory, but the second branch is no longer $p\leftrightarrow q$ interchangeable with it.
@@
Final remarks
While our critical indices are as for the perturbative branch in the usual Liouville theory, the domain of applicability is now broader complimenting applications of the Liouville action in Condensed Matter Theory inspired by [23]. A hint that conformal symmetry of the Nambu-Goto string and the critical Ising model on a dynamical lattice may both be described by the (4,3) minimal model sounds fascinating and deserving future studies.
Equation (4) generically represents a “massive” generalization of the Liouville action. I expect it may have wide applications and help to make an insight in quantum gravity as the usual Liouville action did [24, 25].
The reason why in Eq. (4) I have dropped the terms of higher order in $\tau$, which has resulted in the four-derivative action (10), is twofold. First, this is a simplest extension of the Liouville action, demonstrating how the higher-derivative terms revive. Second, there could be a kind of universality of the higher terms like it happens for the Polyakov string where they are marginal and do not change [12] the results of the usual Liouville action.
I can illustrate this by showing how the nonlocal term (14b) emerges for the Nambu-Goto string in flat space where it precisely comes from averaging over $X^{\mu}$ of EMT for the Nambu-Goto action (2)
$T_{zz}=2\pi\left\langle\lambda^{z{\bar{z}}}\partial X\cdot\partial X+\lambda^{{\bar{z}}{\bar{z}}}\partial X\cdot\bar{\partial}X\right\rangle$ (38)
thanks to the interaction $\lambda^{zz}\partial X\cdot\partial X$ in (2). The emerging $T_{zz}$ is represented by the diagrams
Figure 1: Diagrams representing the average of (38) over $X^{\mu}$. The solid line represents $X^{\mu}$ and the wavy lines represent $\lambda^{ab}$.
[…]
which may be thought as obtained from a six-derivative action. Nonlocal terms emerging from the diagrams in Fig. 1d) etc. will not contribute at one loop. The expectation is that the sum of the contributions of (14b) and (39) to $\langle\hat{\delta}_{\xi}T_{zz}\rangle$ with the singular products computed for the six-derivative action may give the same result as (23) computed for the four-derivative action, like it was the origin of the universality of higher terms for the Polyakov string [12]. I hope to return to this issue elsewhere.
Acknowledgements.
I thank J. Ambjørn, A. Barvinsky, K. Dasgupta, V. Kazakov, C. Kristjansen, A. Morozov, G. Semenoff, M. Shaposhnikov, A. Tseytlin, D. Vasiliev, M. Vasiliev, A. Zamolodchikov, K. Zarembo for useful comments. This work was carried out within the research program of NRC “Kurchatov Institute”.